OpenStudy (anonymous):

Given a circle of radius 18.00 cm. An angle intercepts an arc of 24.0 cm. Determine the sector angle in radians, to the nearest hundredth

OpenStudy (campbell_st):

the arc length is found using the formula \[l = r \theta\] theta is the angle subtended, r = radius and l is the arc length then \[24 = 18\theta\] \[\theta = 24/18 = 4/3 radians\]

OpenStudy (anonymous):

so 1.33

OpenStudy (campbell_st):

oops then its 1.33 radians to the nearest hundredth